1.3) (15 points) Exercise 2.1.18. Except there are 106 parts and the numbers in the first column change from 70, 12 to 81, 11 respectively. Discs of polycarbonate plastic from a supplier are analyzed for scratch and shock resistance. The results from 106 discs are summarized as follows: Shock Resistance High 81 Low Scratch High Resistance Low Let A denote the event that a sample has high shock resistance, and let B denote the event that a disc has high scratch resistance. If a disc is selected at random, determine the number of samples in A nB, A and A UB. stho nasswerds contain
Answer #1

a)

A - high shock resistance

B- high scratch resistance

n(A ad B) = 81

n (A') = 9+5=14

n(A or B) = 81 + 9 + 11 = 101

Please rate